I'm confused on these problems can you show me step by step how to do this?
Janet is mixing 15% glucose solution with a 35% glucose solution. This mixture produces 35 liters if a 19% glucose solution. How many liters of each solution is Janet using in the mixture?

Answer by stanbon(75887)

Janet is mixing 15% glucose solution with a 35% glucose solution.
This mixture produces 35 liters of a 19% glucose solution.
How many liters of each solution is Janet using in the mixture?
-----------------------------------------
Equation:
glucose + glocose = glucose
0.15x + 0.35(35-x) = 0.19*35
Multiply thru by 100 to get:
15x + 35*35 - 35x = 19*35
-20x = -16*35
x = 28 liters (amount of 15% glucose needed in the mixture)
35-x = 35-28 = 5 liters (amount of 35% glucose needed in the mixture)
=======================================================================
Cheers,
Stan H.